Coulomb's Law: Two tiny particles having

charges of +5.00 μC and +7.00 μC are

placed along the x-axis. The +5.00-μC

particle is at x = 0.00 cm, and the other

particle is at x = 100.00 cm. Where on the x-axis

must a third charged particle be placed

so that it does not experience any net

electrostatic force due to the other two

particles?
Answer #1

Charge on the first particle = q1 = 5 C = 5 x 10-6 C

Charge on the second particle = q2 = 7 C = 7 x 10-6 C

Charge on the third particle = q3

Let the third particle be placed on the x-axis at x = d.

Distance of the third particle from the first particle = d

Distance of the third particle from the second particle = 100 - d

Electrostatic force on the third particle due to the first particle = F1 = kq1q3/d2

Electrostatic force on the third particle due to the second particle = F2 = kq2q3/(100-d)2

The third particle experiences zero net force,

F1 = F2

kq1q3/d2 = kq2q3/(100-d)2

q1/d2 = q2/(100-d)2

(5x10-6)/d2 = (7x10-6)/(100-d)2

5/d2 = 7/(100-d)2

5(100-d)2 = 7d2

50000 - 1000d + 5d2 = 7d2

2d2 + 1000d - 50000 = 0

d2 + 500d - 25000 = 0

d = 45.805 or -545.805

d is not negative.

d = 45.805 cm

The third particle should be placed at x = 45.805 cm on the x-axis so that it does not experience any net force due to the other two particles.